# Towards Long-Term User Welfare in Recommender Systems via Creator-Oriented Information Revelation

### v1

@@ 5.2. Effectiveness of LoRe (RQ1)
To evaluate the effectiveness of LoRe in enhancing long-term user welfare, we compare its performance against several baselines on two recommendation datasets. In this experimental setting, the number of creators is fixed at 50 and the number of users at 100. The platform employs DIN (Zhou et al., 2018) as the based recommendation model, and both user and creator agents are powered by LLaMA3-8B (Touvron et al., 2023). In addition, we initialize creators’ trust degrees using a truncated Gaussian distribution, reflecting real-world scenarios where agents exhibit moderate trust levels, and extremely high or low trust is uncommon (Teacy, 2006; Teacy et al., 2012). Denoted by $\mathcal{N}(\mu,\sigma;0,1)$ a truncated Gaussian distribution with mean $\mu$, variance $\sigma$, lower bound 0, and upper bound $1$. Then, each creator’s trust degree is randomly sampled from distribution $\mathcal{N}(0.5,1.0;0,1)$.
Figure 2 shows the changes in cumulative user clicks per round as the number of rounds increases. LoRe consistently achieves higher user welfare than all baselines on both the YouTube and Amazon datasets. Compared with approaches that only adjust recommendation algorithms or creator rewards (e.g., PMMF, FairRec, Reweight), LoRe performs substantially better, indicating that influencing creator behavior through selective information revelation is more effective for improving long-term user welfare.
Existing methods also show inconsistent performance across datasets. For instance, Reweight outperforms all other baselines on YouTube but performs worse than naive information revelation strategies on Amazon. This is primarily due to the more skewed user distribution in the Amazon dataset compared to the YouTube dataset, i.e., the majority of users in the Amazon dataset favor a particular genre of content, while fewer users favor other genres, which means that existing methods may significantly degrade in performance under such conditions.
Naive information revelation strategies (Most History Click, Most Click) also display unstable performance: they outperform w/o Signal on Amazon but are similar to it on YouTube. This suggests that naive strategies lack robustness across recommendation environments. In contrast, LoRe consistently achieves the best long-term welfare, demonstrating that strategic information revelation is not only more effective than naive approaches but also more robust across different environments.
Figure 4. Training time cost and the influence of trust degree.
@@ 5.3. Inherent insight of LoRe (RQ2)
To better understand the inherent insight of LoRe, we analyze the dynamics of RS from both macro and micro perspectives. At the macro level, we examine the content diversity and the average number of active creators during the interactions among the platform, creators, and users. At the micro level, we investigate the variation in the number of content genres produced by individual creators under different methods.
@@
In our simulations, we assumed that a creator would exit after receiving no clicks for 10 consecutive rounds. We calculated the average number of active creators under different methods across datasets. As shown in Table 2, our method retains more creators than other baselines. This improvement arises because the platform can influence creators’ behavior by suggesting them creating specific content genre, thereby guiding them to produce content that attracts more clicks.
Micro perspective. We further analyze the variation in the number of content genres produced by each creator. As shown in Figure 3, the x-axis represents the creator ID, and the y-axis indicates the number of distinct genres created by each individual. Obviously, compared to the baseline "w/o Signal", our approach leads to an increase in the number of genres per creator, which directly contributes to the overall improvement in content diversity.
@@ 5.4. Robustness Analysis (RQ3)
Training cost. We evaluate the convergence rounds and net training time of our algorithm at different scales to demonstrate its scalability in large-scale simulations. As shown in Figure 4 (a), our algorithm requires only a few rounds to converge. When the number of agents is 50, it converges in fewer than 70 rounds; when the number of agents increases to 1,000, convergence is achieved in fewer than 40 rounds. In addition, the net training time per round remains within an acceptable range. Even with 1,000 agents, the net training time per round is less than 4 seconds. These results confirm that our method scales effectively to large-scale, real-world scenarios.
@@
Influence of creator simulation models. As discussed in Section 3.3, our method is capable of handling various creator simulation models. To verify this and evaluate its robustness, we conduct experiments across different creator simulation models. Specifically, we consider both LLM-based simulation models: CreAgent (Ye et al., 2025) (with Mistral-7B and Qwen2.5-7B as backbones), and heuristic-based models: Creator Feature Dynamics (CFD) (Lin et al., 2024; Zhan et al., 2021), where creators adjust their creation strategies using user feedback as gradients scaled by a learning rate, and SimuLine (Zhang et al., 2023), where creators’ subsequent creations are determined by probabilistic sampling based on the number of clicks in previous steps.
@@
Figure 5 presents the performance comparison between our method and the baselines under different creator simulation models. The results clearly show that, regardless of whether the simulation model is LLM-based or heuristic-based, our approach consistently outperforms all baselines. In contrast, existing recommendation-algorithm-based methods (i.e., P-MMF and FairRec) exhibit inconsistent performance across different creator simulation models. For instance, FairRec performs best among baselines under the Mistral-7B, Qwen2.5-7B, and SimuLine simulation models but is outperformed by P-MMF under the CFD model. Moreover, Reweight, which influences creator behavior by adjusting their rewards, performs relatively poorly. This is because it is designed for a specific creator decision model and thus may fail to generalize well when the decision model changes.
@@
Table 3. The average user clicks achieved through combining LoRe with FairRec or P-MMF on YouTube and Amazon datasets.
@@
LoRe (ours) 137.4954 95.1499
P-MMF 108.5688 51.2752
FairRec 113.2789 46.1865
LoRe + P-MMF 141.2263 101.9969
LoRe + FairRec 146.0092 98.6177
Influence of creators’ trust degree. Although platforms can provide various suggestions to creators, real-world creators differ in their degree of trust toward the platform. Highly trusting creators are more likely to follow the platform’s suggestions, whereas those with low trust are less likely to be influenced. In addition, in practice, creators’ trust levels may evolve over time according to the number of clicks they receive Prasad et al. (2023).
To evaluate the effectiveness of our method under different levels of creator trust, we conduct experiments where creators’ trust degrees are sampled from distributions with varying expectations. Specifically, we conduct experiments with $\mu\in\{0.4,0.6,0.8\}$ and $\sigma=1.0$. Moreover, we assume that creators’ trust degrees change dynamically over time. Let $d_{i}^{t}$ denote creator $i$’s trust degree in round $t$ and $r_{i}^{t}$ represent the number of clicks received by creator $i$ in round $t$. Then, creator $i$’s trust in round $t+1$ is updated as $d_{i}^{t+1}=d_{i}^{t}+\frac{r_{i}^{t+1}-r_{i}^{t}}{r_{i}^{t}}$, if they follow the platform’s suggestion. This means that a creator’s trust increases if following the platform’s suggestion yields more clicks, and decreases otherwise.
Figure 4 (b) shows the user clicks under trust degree distributions with different expectations on the YouTube dataset. We compare our method with Reweight, which performs best among the baselines on YouTube. The results show that although our method’s performance fluctuates slightly across different trust distributions, it consistently outperforms Reweight, demonstrating robustness to varying levels of creator trust. Furthermore, while performance of our method decreases slightly as the expected trust degree decreases, the decline is marginal. This trend is reasonable, as lower trust leads to fewer creators adopting the platform’s suggestions. These findings highlight the importance of platforms maintaining their credibility by providing beneficial suggestions to creators.
Incorporated with existing methods. We integrate our method with existing approaches that focus on recommendation algorithm design, such as P-MMF and FairRec. We exclude Reweight from this combination, as both LoRe and Reweight aim to enhance user welfare by influencing creators’ behavior. Table 3 presents the average user clicks achieved by combining our method with other approaches on the YouTube and Amazon datasets. It is evident that combining our method with these approaches further improves the performance. Moreover, LoRe + FairRec achieves better performance on the YouTube dataset, whereas LoRe + P-MMF performs better on the Amazon dataset. These results demonstrates that our approach is compatible with other methods and highlights its scalability.
@@ 6. Conclusions
In this paper, we reveal that influencing creators’ behaviors could be a better way to improve user welfare. Inspired by the idea of Bayesian persuasion, we model the platform as the sender and the creators as the receivers. To overcome the problem of bounded rationality assumption, we model the creators as bounded rational agents. Then we propose a novel method, LoRe, which aims to learn how to influence creators by revealing information to them. Lore introduces a reinforcement learning algorithm to address the limitations imposed by the strong assumption in Bayesian persuasion. Extensive experiments on two real-world datasets demonstrate the effectiveness of our approach and show that it can be seamlessly integrated into the existing RS ecosystem, as well as combined with other methods that focus on recommendation algorithm design.
Appendix
Appendix A Examples of Content Suggestion in Real Applications
Figure 6. Examples of creation suggestions in real-world applications across different short video platforms, each of which has hundreds of millions of users.
As the number of creators on short-video platforms continues to grow, providing creation suggestions to guide creators, thereby indirectly improving user experience, has become a key component of platform operations. Figure 6 presents examples of creation suggestions on short-video content platforms, each of which typically serves hundreds of millions of users.
Although current commercial platforms have begun to adopt creation suggestions, they are typically limited to simple suggestions, such as suggesting visual effects or the most trending topics. Effectively guiding creators in real-world industrial settings remains an open challenge. To bridge this gap, we aims to explore how to design more effective suggestion strategies in real-world platform environments to enhance overall user welfare.
Appendix B Omitted Definitions in Section 4
We respectively define the state value function $V^{\pi}(\bm{s}^{t})$, the state-action value function $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$as:
$V^{\pi}(\bm{s}^{t})=\operatorname{\mathbb{E}}_{(\bm{a}^{t},\bm{s}^{t+1},\dots)\sim\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right],$
$Q^{\pi}(\bm{s}^{t},\bm{a}^{t})=\operatorname{\mathbb{E}}_{(\bm{s}^{t+1},\bm{a}^{t+1},\dots)\sim\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right].$
Appendix C More Detailed Experimental Setups
Our experimental server is equipped with 28 Intel(R) Core(TM) i9-10940X CPUs, 128GB of RAM, and 4 NVIDIA RTX 4090 GPUs, providing a total of 96 GB of VRAM.
Appendix D Additional Experiments
D.1. Influence of Different Recommendation Models
Figure 7. Comparison of user welfare with and without our method under the RS environment with two different recommendation models, i.e, MF and BPR, respectively.
In real-world RS environments, platforms adopt different recommendation algorithms to select content from the item pool for users. To verify the effectiveness of our method under various recommendation models, we conduct experiments using different recommendation algorithms as the platform’s policy. We select two of the most widely used recommendation models: Matrix Factorization (MF) (Koren et al., 2009) and Bayesian Personalized Ranking (BPR) (Rendle et al., 2012).
As illustrated in Figure 7, across different recommendation models, our method consistently achieves better user welfare compared to the w/o signal baseline. Specifically, when using MF as the recommendation model, the user welfare with information signaling is 1.18 times that of the no-signal case (at the last round of simulation). When using BPR, this number increases to 1.26 times.
Appendix E Limitations
In our work, we use large language models (LLMs) to simulate the behaviors of users and content creators. Although previous work has verified the effectiveness of the LLM-empowered simulation agents in simulating users click behavior and creator content creation behavior, there is limited research on whether such agents can effectively simulate creators’ responses to platform-provided suggestions. Therefore, how to simulate the behavior of real-world creators after receiving platform suggestions is also a key challenge and represents a promising direction for future research.
Appendix F Social Impacts
Our method can effectively enhance the long-term user welfare in RS. This not only improves user satisfaction and engagement over time but also contributes to the development of a healthier and more sustainable ecosystem for content consumption and creation. To the best of our knowledge, our method does not pose any adverse effects on society.

### v2

@@ 5.2. Effectiveness of LoRe (RQ1)
To evaluate the effectiveness of LoRe in enhancing long-term user welfare, we compare its performance against several baselines on two recommendation datasets. In this experimental setting, the number of creators is fixed at 50 and the number of users at 100. The platform employs DIN (Zhou et al., 2018) as the based recommendation model, and both user and creator agents are powered by LLaMA3-8B (Touvron et al., 2023).
Figure 2 illustrates the evolution of user clicks per round as the number of interaction rounds increases, while Table 2 reports the average number of user clicks over $T$ rounds for all methods and datasets. Across both the YouTube and Amazon datasets, LoRe consistently achieves higher user welfare than all baselines. Compared with methods that focus solely on adjusting recommendation algorithms or creator rewards (e.g., PMMF, FairRec, and Reweight), LoRe exhibits substantially better performance, indicating that influencing creators’ creation behavior through information revelation is a more effective mechanism for improving long-term user welfare.
Existing baselines also exhibit inconsistent performance across datasets. For example, Reweight outperforms other baselines on the YouTube dataset but performs worse than naive information revelation strategies on the Amazon dataset. This discrepancy is primarily attributable to the more skewed user preference distribution in the Amazon dataset, where a large proportion of users favor a single content genre, while relatively few users prefer other genres. Under such conditions, the performance of existing methods can degrade significantly.
Naive information revelation strategies (Most History Click and Most Click) likewise demonstrate unstable performance: they outperform the w/o Signal baseline on the Amazon dataset but perform comparably on the YouTube dataset. This further indicates that naive strategies lack robustness across different recommendation environments. In contrast, LoRe consistently achieves the highest long-term welfare across all settings, demonstrating that carefully designed information revelation is not only more effective than naive approaches but also more robust across diverse environments.
@@ 5.3. Inherent insight of LoRe (RQ2)
To better understand the inherent insight of LoRe, we analyze the dynamics of RS from both macro and micro perspectives. At the macro level, we examine the content diversity and the average number of active creators during the interactions among the platform, creators, and users. At the micro level, we investigate how the number of item genres produced by individual creators varies across different methods.
Figure 4. Training time cost and the influence of trust degree. The left figure presents the number of convergence rounds and training time per rounds for LoRe under different agent number. The right figure shows the user clicks of LoRe under different trust degree.
@@
In our simulations, a creator would exit after receiving no clicks for 10 consecutive rounds. We calculated the average number of active creators under different methods across datasets. As shown in Table 2, our method retains more creators than other baselines. This improvement arises because the platform can influence creators’ behavior by suggesting them creating specific content genre, thereby guiding them to produce content that attracts more clicks.
Micro perspective. We further analyze the variation in the number of content genres produced by each creator. As shown in Figure 3, the x-axis represents the creator ID, and the y-axis indicates the number of distinct genres created by each individual. Obviously, compared to the baseline w/o Signal, our approach leads to an increase in the number of genres per creator, which directly contributes to the overall improvement in content diversity.
@@ 5.4. Robustness Analysis (RQ3)
Training cost. We evaluate the convergence rounds and net training time of our algorithm at different scales to demonstrate its scalability in large-scale simulations. As shown in Figure 4 (a), our algorithm requires only a few rounds to converge. When the number of agents is 50, it converges in fewer than 70 rounds; when the number of agents increases to 1,000, convergence is achieved in fewer than 40 rounds. In addition, the net training time per round remains within an acceptable range. Even with 1,000 agents, the net training time per round is less than 4 seconds. These results confirm that our method scales effectively to large-scale scenarios.
@@
Influence of creator simulation models. To evaluate our method’s robustness, we conduct experiments under multiple creator simulation models. Specifically, we consider both LLM-based simulation models: CreAgent (Ye et al., 2025) (with Mistral-7B and Qwen2.5-7B as backbones), and heuristic-based models: Creator Feature Dynamics (CFD) (Lin et al., 2024; Zhan et al., 2021), in which creators adjust their creation behavior using user feedback as gradients scaled by a learning rate, and SimuLine (Zhang et al., 2023), where creators’ creations are determined by probabilistic sampling based on the number of clicks in previous steps.
When LLM-based simulation models are used, the platform’s suggestions can be directly incorporated into creators’ prompts, thereby influencing their behavior. However, this mechanism is not applicable to heuristic-based models. To simulate the effect of platform suggestions in such settings, inspired by  Prasad et al. (2023), we design a trust-based model. Specifically, each creator $c_{i}$ is assigned a trust degree $d_{i}\in[0,1]$, representing the probability that the creator follows the platform’s suggestion upon receiving it. In our experiments, creators’ trust degrees are initialized using a truncated Gaussian distribution to reflect real-world scenarios in which agents typically exhibit moderate trust, while extremely high or low trust levels are uncommon (Teacy, 2006; Teacy et al., 2012). Denoted by $\mathcal{N}(\mu,\sigma;0,1)$ a truncated Gaussian distribution with mean $\mu$, variance $\sigma$, lower bound 0, and upper bound $1$. Then, each creator’s trust degree is randomly sampled from distribution $\mathcal{N}(0.5,1.0;0,1)$.
@@
Figure 5 reports the performance comparison between our method and the baselines under different creator simulation models. The results show that, regardless of whether the simulation model is LLM-based or heuristic-based, our approach consistently outperforms all baselines. In contrast, existing recommendation-algorithm-based methods (i.e., P-MMF and FairRec) exhibit inconsistent performance across different creator simulation models. For instance, FairRec performs best among baselines under the Mistral-7B, Qwen2.5-7B, and SimuLine simulation models but is outperformed by P-MMF under the CFD model. Moreover, Reweight, which influences creator behavior by adjusting their rewards, performs relatively poorly. This is because it is designed for a specific creator decision model and thus may fail to generalize well when the decision model changes.
Influence of creators’ trust degree. When heuristic-based models are used to simulate creators’ behavior, we assume that each creator has a degree of trust on the platform. In real-world RS, creators naturally differ in their trust. Highly trusting creators are more likely to follow platform suggestions, whereas creators with low trust are less susceptible to such influence. Moreover, creators’ trust may evolve over time as a function of the feedback they receive(Prasad et al., 2023). To evaluate the effectiveness of our method under different levels of creator trust, we conduct experiments where creators’ initial trust degrees are sampled from distributions with different expectations. Specifically, we conduct experiments with $\mu\in\{0.4,0.6,0.8\}$ and $\sigma=1.0$. Moreover, we assume that creators’ trust degrees change dynamically over time. Let $d_{i}^{t}$ denote creator $i$’s trust degree in round $t$ and $r_{i}^{t}$ represent the number of clicks received by creator $i$ in round $t$. Then, creator $i$’s trust in round $t+1$ is updated as $d_{i}^{t+1}=d_{i}^{t}+\frac{r_{i}^{t+1}-r_{i}^{t}}{r_{i}^{t}}$, if they follow the platform’s suggestion. This means that a creator’s trust increases if following the platform’s suggestion yields more clicks, and decreases otherwise.
Figure 4 (b) shows the user clicks under trust degree distributions with different expectations on the YouTube dataset. We compare our method with Reweight, since it performs best among the baselines on YouTube. The results show that although our method’s performance fluctuates slightly across different trust distributions, it consistently outperforms Reweight, demonstrating robustness to varying levels of creator trust. Furthermore, while performance of our method decreases slightly as the expected trust degree decreases, the decline is marginal. Overall, these findings underscore the importance of maintaining platform credibility by providing consistently beneficial suggestions to creators.
Figure 6. Comparison of user welfare with our method and baselines under the RS environment with two different recommendation models, i.e, MF and BPR, respectively.
Influence of different recommendation algorithms. In real-world RS environments, platforms employ different recommendation algorithms to select content from the item pool for users. To verify the effectiveness of our method under various recommendation models, we conduct experiments using different recommendation algorithms as the platform’s recommendation policy. Specifically, we select two of the most widely used recommendation models: Matrix Factorization (MF) (Koren et al., 2009) and Bayesian Personalized Ranking (BPR) (Rendle et al., 2012). As illustrated in Figure 6, our method consistently achieves higher user welfare than the baselines across different recommendation algorithms. This result demonstrates that the effectiveness of LoRe is not tied to a specific recommendation model, but instead generalizes well across different algorithmic.
@@
Table 3. The average user clicks achieved through combining LoRe with FairRec or P-MMF on different datasets.
@@
P-MMF $108.569\pm 1.213$ $51.275\pm 0.984$
FairRec $113.279\pm 1.146$ $46.187\pm 0.813$
LoRe (ours) $137.495\pm 2.523$ $95.150\pm 0.687$
LoRe + P-MMF $\underline{141.226}\pm 2.266$ $\bm{101.997}\pm 1.084$
LoRe + FairRec $\bm{146.009}\pm 1.742$ $\underline{98.618}\pm 0.994$
Incorporated with existing methods. We integrate our method with existing approaches that focus on recommendation algorithm design, including P-MMF and FairRec. We exclude Reweight from this combination, as both LoRe and Reweight aim to enhance user welfare by influencing creators’ creation behavior. Table 3 presents the average number of user clicks achieved by combining our method with these approaches on the YouTube and Amazon datasets. The results show that integrating LoRe with these methods further improves performance across both datasets. In particular, LoRe + FairRec achieves higher user welfare on the YouTube dataset, whereas LoRe + P-MMF performs better on the Amazon dataset. These findings demonstrate that LoRe is compatible with existing recommendation algorithm and highlight its scalability.
6. Social Impacts
Our method effectively enhances long-term user welfare in recommendation systems. Beyond improving user satisfaction and engagement over time, it also contributes to the development of a healthier and more sustainable ecosystem for content creation and consumption. A potential concern is that platform might misuse such suggestion strategy to manipulate creators’ behavior, benefiting some creators at the expense of others. However, the proposed suggestion strategy does not compel creators’ actions, that is creators retain full autonomy over the content they create. Moreover, if the platform were to provide suggestions that systematically harm creators’ earnings, it would erode the platform’s credibility, reduce creators’ trust, and ultimately diminish user welfare. Thus, platforms lack incentives to issue harmful or manipulative suggestions.
@@ 7. Conclusions
In this paper, we demonstrate that influencing creators’ creation behavior could be a better way to improve long-term user welfare compared to methods relying on recommendation algorithm design. Inspired by the idea of cheap talk, we model the platform as the sender and creators as the receivers. To address the challenges of applying cheap talk to RS, we propose a novel method, LoRe, which formulates the dynamic interaction between the platform and creators as a MDP. LoRe employs a reinforcement learning algorithm to optimize the platform’s suggestion strategy. Extensive experiments on two real-world datasets demonstrate the effectiveness of our approach and show that it can be seamlessly integrated into the existing RS ecosystem, as well as combined with other methods that focus on recommendation algorithm design.